CH/Pubs/01

Publication List

Cumulative: All Categories Author: N. Curtis and W.N. Catford Date: 31st October 2001

Contents

1. List of Doctoral Theses Produced

2. List of Publications in Refereed Journals 3. Papers Accepted for Publication

4. Papers Recently Submitted

5. Papers in Preparation

6. Invited Talks at Conferences

7. Conference Contributions

8. Other Output

1 List of Doctoral Theses Produced

1. “Resonant particle spectroscopy”, S.C. Allcock, University of Oxford, 1986.

2. “Spin and structure of ^{12}C+^{12}C resonances”, P.R. Keeling, University of Oxford, 1987.

3. “Symmetric and near-symmetric breakup of sd-shell nuclei”, S.J. Bennett, University of Birmingham, 1989.

4. “Measurement of the spins of symmetrically fissioning states in ^{24}Mg”, M. Freer, University of Birmingham, 1991.

5. “Breakup studies with a ^{23}Na beam”, N.S. Jarvis, University of York, 1991.

6. “Symmetric fissioning states at high excitation in ^{24}Mg”, J.T. Murgatroyd, University of Birmingham, 1992.

7. “Symmetric breakup studies in ^{24}Mg”, C.D. Jones, University of York, 1993.

8. “High excitation breakup states in ^{24}Mg and related detector development”, M.J. Leddy, University of Birmingham, 1994.

9. “^{16}O+^{8}Be breakup of ^{24}Mg”, J.S. Pople, University of Birmingham, 1994.

10. “A study of highly deformed α–cluster structures in light nuclei”, P.M. Simmons, University of Oxford, 1995.

11. “Study of super-deformed α–cluster states in light nuclei,” J. Zhang, University of Oxford, 1995.

12. “The ^{12}C+^{12}C ‘6–α Chain State’ Resonance,”

S.P.G. Chappell, University of York, 1995.

13. “^{12}C and ^{16}O Breakup of ^{24}Mg, ^{28}Si and ^{32}S,”

N. Curtis, University of Birmingham, 1995.

14. “A study of the ^{12}C(^{16}O,^{12}C(E_{x}=7.65 MeV))^{16}O reaction,” S.P. Fox, University of York, 1995.

15. “A search for ^{16}O + ^{16}O + ^{16}O breakup of ^{48}Cr ”

A.St.J. Murphy, University of Birmingham, 1996.

16. “Symmetric fission of ^{24}Mg and ^{56}Ni,”

S.M. Singer, University of Surrey, 1996.

17. “A search for shape isomers in ^{28}Si, ^{24}Mg and ^{20}Ne linked by successive alpha decay,” G.R. Kelly, University of Birmingham, 1996.

18. “The ^{16}O + ^{8}Be exit channel in the vicinity of the ^{12}C* + ^{12}C* E_{cm}=32.5 MeV resonance,” P. Lee, University of York, 1997.

19. “Investigation of the ^{12}C + ^{12}C E_{cm}=32.5 MeV resonance,”

R.A. Le Marechal, University of Birmingham, 1997.

20. “A Multi-Element Gas-Hybrid Array for charged particle spectroscopy,” R.L. Cowin, University of York, 1997.

21. “Search for a shape eigenstate in ^{32}S,”

S.J. Hoad, University of Birmingham, 1998.

22. “Couple channels calculations for the scattering of deformed light nuclei,” P.E. Fry, University of Oxford, 1998.

23. “Investigation of resonance phenomena in the ^{16}O + ^{16}O system,”

G.K. Dillon, University of York, 1999.

24. “Breakup reactions populating cluster states in ^{28}Si and ^{24}Mg,”

M. Shawcross, University of Surrey, 1999.

25. “An angular correlation technique for the investigation of nuclear reaction mechanisms with an application to the ^{12}C(^{24}Mg,^{12}C ^{12}C)^{12}C reaction,” C.D. Freeman, University of Birmingham, 1999.

26. “Couple channels calculations for the scattering of deformed light heavy ions: A challenge to the standard approach,” I. Boztosun, University of Oxford, 2000.

27. “Probing nuclear molecular analogue states in carbon, boron and beryllium isotopes,” P.J. Leask, University of Birmingham, 2000.

28. “New high spin resonances in ^{12}C + ^{12}C scattering,”

C.A. Bremner, University of Oxford, 2000.

29. “The unbound nucleus ^{13}Be,”

K.L. Jones, University of Surrey, 2000.

30. “The effect of breakup in the ^{9}Be + ^{208}Pb system,”

R.J. Woolliscroft, University of Birmingham, 2001.

2 List of Publications in Refereed Journals

2.1 1984

1. “Reaction mechanism and spin information in the sequential breakup of ^{18}O at 82 MeV”, W.D.M. Rae and R.K. Bhowmik, Nucl. Phys. A420 (1984) 320–350

2. “Spin measurements and reaction mechanism in the inelastic scattering of ^{18}O on 12C”, R.K. Bhowmik, W.D.M. Rae and B.R. Fulton, Phys. Lett. 136B (1984) 149–152

3. “Excitation of Anomalous α-^{14}C Resonances in the Inelastic Scattering of ^{18}O on a ^{12}C Target,” W.D.M. Rae and R.K. Bhowmik, Nucl. Phys. A427 (1984) 142

2.2 1985

1. “Angular correlations in heavy ion reactions”, S. Marsh and W.D.M. Rae, Phys. Lett. 153B (1985) 21

2. “High spin α-cluster resonances in ^{16}O,”

W.D.M. Rae, S.C. Allcock, S. Marsh and B.R. Fulton, Phys. Lett. 156B (1985) 167–171

3. “A study of ^{19}F by inelastic scattering on a ^{12}C target at 82 MeV and 144 MeV using the techniques of resonant particle spectroscopy,” A.E. Smith, S.C. Allcock, W.D.M. Rae, B.R. Fulton and D.W. Banes, Nucl. Phys. A441 (1985) 701

4. “High-spin α-cluster Resonances in ^{16}O,”

W.D.M. Rae, S.C. Allcock, S. Marsh and B.R. Fulton, Phys. Lett. 156B (1985) 167

5. “Stretching, Anti-stretching and Particle Alignment in Rapidly Rotating Cluster Nuclei and Nuclear Quasi-molecules,” W.D.M. Rae and S. Marsh, Phys. Lett.161B (1985) 251

6. “Search for heavy-ion resonances in the reactions^{12}C(^{20}Ne,α) and^{12}C(^{20}Ne,^{8}Be),” N.R. Fletcher, W.D.M. Rae, A.E. Smith, S.C. Allcock and D.L. Gay, Phys. Rev. C32 (1985) 1911

2.3 1986

1. “Symmetric Fission of ^{24}Mg following Inelastic Scattering,”

B.R. Fulton, S.J. Bennett, C.A. Ogilvie, J.S. Lilley, D.W. Banes, W.D.M. Rae, S.C. Allcock, R.R. Betts, A.E. Smith, Phys. Lett. 181B (1986) 233

2. “The structure of ^{24}Mg using the cranked cluster model”, S. Marsh and W.D.M. Rae, Phys. Lett. 180B (1986) 185

3. “Symmetric fission of ^{24}Mg following inelastic scattering,”

B.R. Fulton, S.J. Bennett, C.A. Ogilvie, J.S. Lilley, D.W. Banes, W.D.M. Rae, S.C. Allcock, R.R. Betts and A.E. Smith, Phys. Lett. 181B (1986) 233

2.4 1987

1. “Spins of ^{12}C + ^{12}C resonances via the ^{12}C(^{12}C,αα)^{16}O reaction,”

W.D.M. Rae, P.R. Keeling and S.C. Allcock, Phys. Lett. 184B (1987) 133–138

2. “The Gross Structure Resonance Near E_{CM} = 25 MeV in ^{12}C + ^{12}C scattering,” W.D.M. Rae, P.R. Keeling and A.E. Smith, Phys. Lett. 198B (1987) 49

3. “10^{+} states in ^{20}Ne,”

S.C. Allcock, W.D.M. Rae, P.R. Keeling, A.E. Smith, B.R. Fulton and D.W. Banes, Phys. Lett. 201B (1987) 201

2.5 1988

1. “Clustering phenomena and shell effects in nuclear structure and reactions,” W.D.M. Rae, Int. J. Mod. Phys. A3 (1988) 1343–1372

2.6 1989

1. ^{16}O+^{8}Be cluster structure in ^{24}Mg,”

B.R. Fulton, S.J. Bennett, M. Freer, R.D. Page, P.J. Woods, S.C. Allcock, A.E. Smith, W.D.M. Rae and J.S. Lilley, J. Phys. Soc. Jpn. 58 (1989) Suppl. p.590

2. “^{16}O+^{8}Be breakup of ^{24}Mg,”

B.R. Fulton, S.J. Bennett, M. Freer, R.D. Page, P.J. Woods, S.C. Allcock, A.E. Smith, W.D.M. Rae and J.S. Lilley, Phys. Lett. 232B (1989) 56

2.7 1990

1. “A compact high-efficiency ^{8}Be detector,”

S.C. Allcock, P.R. Keeling, W.D.M. Rae, A.E. Smith, S.J. Bennett, M. Freer, B.R. Fulton, R.D. Page and P.J. Woods, Nucl. Instr. Meth. A289 (1990) 213

2. “Fission of light nuclei,”

B.R. Fulton and W.D.M. Rae, J. Phys. G16 (1990) 333–365

2.8 1991

1. “Evidence for excitation of ^{24}Mg resonant states through the ^{12}C(^{16}O,^{12}C^{12}C)^{4}He reaction at 85 MeV,” E. Costanzo, M. Lattuada, S. Romano, D. Vinciguerra, N. Cindro, M. Zadro, M. Freer, B.R. Fulton and W.D.M. Rae, Europhys. Letts. 14 (1991) 221

2. “Search for large cluster breakup in ^{40}Ca,”

B.R. Fulton, S.J. Bennett, M. Freer, J.T. Murgatroyd, S.C. Allcock, W.D.M. Rae, A.E. Smith, B. Back, R.R. Betts, S. Sanders and F. Videbaek, J. Phys. G17 (1991) 555–559

3. “Excitation of ^{24}Mg states through interaction of 85 MeV ^{16}O ions with ^{12}C and 24Mg targets,” E. Costanzo, M. Lattuada, S. Romano, D. Vinciguerra, N. Cindro, M. Zadro, M. Freer, B.R. Fulton and W.D.M. Rae, Phys. Rev. C44 (1991) 111

4. “Near symmetric breakup of ^{25}Mg,”

G.J. Gyapong, N.S. Jarvis, D.L. Watson, S.J. Bennett, M. Freer, B.R. Fulton, J.T. Murgatroyd, R.A. Hunt, W.D.M. Rae and A.E. Smith, Phys. Rev. C44 (1991) 520

5. “Spin measurement for symmetric fission states of ^{24}Mg — a new angle on the 12C+^{12}C interaction,” B.R. Fulton, S.J. Bennett, M. Freer, J.T. Murgatroyd, G.J. Gyapong, N.S. Jarvis, C.D. Jones, D.L. Watson, J.D. Brown, W.D.M. Rae, A.E. Smith and J.S. Lilley, Phys. Lett. 267B (1991) 325

6. “Symmetric and near symmetric fission of ^{24}Mg, ^{28}Si and ^{32}S,”

S.J. Bennett, M. Freer, B.R. Fulton, J.T. Murgatroyd, P.J. Woods, S.C. Allcock, W.D.M. Rae, A.E. Smith, J.S. Lilley and R.R. Betts, Nucl. Phys. A534 (1991) 445–460

2.9 1992

1. “Rise time effects in position sensitive detectors,”

S.J. Bennett, M. Freer, B.R. Fulton and J.T. Murgatroyd, Nucl. Instr. Meth. A311 (1992) 105

2. “Hyperdeformed states in ^{36}Ar and ^{48}Cr in the cranked cluster model”, W.D.M. Rae and A.C. Merchant, Phys. Lett. B279 (1992) 207–211

3. “The shape eigenstate: a new kind of resonance”, W.D.M. Rae, A.C. Merchant and B. Buck, Phys. Rev. Lett. 69 (1992) 3709–3712

4. “Alpha chain states in 4N–nuclei from ^{20}Ne to ^{32}S”, A.C. Merchant and W.D.M. Rae, Phys. Rev. C46 (1992) 2096–2097

5. “Systematics of alpha chain states in 4N–nuclei”, A.C. Merchant and W.D.M. Rae, Nucl. Phys. A549 (1992) 431–438

2.10 1993

1. “Experimental techniques for inverse kinematic reactions using thick targets applied to ^{1}H + ^{23}Na → ^{12}C+ ^{12}C,” S.J. Bennett, N.M. Clarke, M. Freer, B.R. Fulton, S.J.A. Hall, O. Karban, J.T. Murgatroyd, G. Tungate, G.J. Gyapong, N.S. Jarvis, D.L. Watson, W.D.M. Rae, A.E. Smith, J.S. Lilley, P.J. Woods and R.D. Page, Nucl. Instr. Meth. A332 (1993) 476

2. “Energy dependence of the angular distributions of a new resonance seen in 12C(^{12}C,^{12}C(0_{2}^{+}))^{12}C(0_{2}^{+})”, A.C. Merchant and W.D.M. Rae, J. Phys. G19 (1993) L89–L93.

3. “Comment on Evidence for superdeformed shape isomeric states in ^{28}Si at excitations above 40 MeV through observations of selective particle decays of ^{16}O + ^{12}C resonances in ^{8}Be and alpha channels”, J.Zhang, A.C.Merchant and W.D.M.Rae, Phys. Rev. C48 (1993) 2117–2119

4. “Shape eigenstates and other one- and twodimensional alpha cluster structures in light nuclei.”, W.D.M.Rae and A.C.Merchant, Mod. Phys. Letts. A8 (1993) 2435–2447 5. “Systematics of 2-dimensional alpha cluster configurations in 4N-nuclei from ^{12}C

to ^{44}Ti”, J.Zhang and W.D.M.Rae, Nucl. Phys. A564 (1993) 252–270

2.11 1994

1. “Spin assignments for states in ^{20}Ne via the ^{12}C(^{16}O, ^{20}Ne^{∗})^{8}Be transfer reaction”, W.D.M.Rae, S.C.Allcock and J.Zhang, Nucl. Phys. A568 (1994) 271–286

2. “Spin assignments for states in ^{16}O via the ^{12}C(^{16}O, ^{16}O^{∗})^{8}Be transfer reaction”, W.D.M.Rae, S.C.Allcock and J.Zhang, Nucl. Phys. A568 (1994) 286–305

3. “Resonances in ^{24}Mg revealed in the ^{12}C(^{20}Ne,^{12}C^{12}C)^{8}Be reaction,”

B.R. Fulton, S.J. Bennett, J.T. Murgatroyd, N.S. Jarvis, D.L. Watson, W.D.M. Rae, Y. Chan, D. DiGregorio, J. Scarpaci, J. Suro Prerez and R.G. Stokstad, J. Phys. G20 (1994) 151–157

4. “Super- and hyperdeformed states in ^{56}Ni”, J.Zhang, A.C.Merchant and W.D.M.Rae, Phys. Rev. C49 (1994) 562–565

5. “Calculations of Particle Decay Widths of Hyperdeformed Nuclei Using Prolate Spheroidal Coordinates”, A.C.Merchant and W.D.M.Rae, Nucl. Phys. A571 (1994) 43–66

6. “Geometric derivations of the second constant of motion for an elliptic ‘billiard’ and other results”, J. Zhang, A.C. Merchant and W. D. M. Rae, Eur. J. Phys. 15 (1994) 133–138

7. “Systematics of some 3–dimensional α–cluster configurations in 4N nuclei from 16O to ^{44}Ti”, J. Zhang, W. D. M. Rae and A.C. Merchant, Nucl. Phys. A575 (1994) 61–71

8. “Geometry and collectivity in the Bloch-Brink α-cluster model”, W.D.M.Rae, A.C.Merchant and J.Zhang, Phys. Letts. B321 (1994) 1–5

9. “Triangular alpha cluster geometries and harmonic oscillator shell structure”, W.D.M.Rae and J.Zhang, Mod. Phys. Letts. A9 (1994) 599–607 10. “Nuclear molecules and nuclear clusters,”

B.R. Fulton, Z. Phys. A349 (1994) 227–232

11. “Alpha chain states in 4N-nuclei”, A.C.Merchant and W.D.M.Rae, Z. Phys. A349 (1994) 243–248

12. “The breakup of ^{24}Mg into ^{16}O+^{8}Be” J.S. Pople, N.M. Clarke, B.R. Fulton, M.J. Leddy, J.T. Murgatroyd, Y. Chan, R.G. Stokstad, W.D.M. Rae, W.N. Catford, S.P. Fox, G.J. Gyapong, D.L. Watson and S.J. Bennett, Z. Phys. A349 (1994) 349–350

13. “Spin and structure of ^{12}C+ ^{12}C resonances I”, W.D.M. Rae and P.R. Keeling, Nucl. Phys. A575 (1994) 175–206

14. “Spin and structure of ^{12}C+ ^{12}C resonances II”, W.D.M. Rae and P.R. Keeling, Nucl. Phys. A575 (1994) 207–220

15. “Reaction mechanism for the symmetric breakup of ^{24}Mg following interaction with ^{12}C”, G.J. Gyapong, D.L. Watson, W.N. Catford, N.M. Clarke, S.J. Bennett, M. Freer, B.R. Fulton, C.D. Jones, M.J. Leddy, J.T. Murgatroyd, W.D.M. Rae and P.M. Simmons, Nucl. Phys. A579 (1994) 207–224

16. “The construction and operation of a hybrid gas-silicon detector for studies of cluster breakup reactions,” N. Curtis, A. St. J. Murphy, M.J. Leddy, J.S. Pople, N.M. Clarke, M. Freer, B.R. Fulton, S.J. Hall, G. Tungate, R.P. Ward, S.M. Singer, W.N. Catford, G.J. Gyapong, R.A. Cunningham, J.S. Lilley, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, P.M. Simmons, R.A. Hunt, A.C. Merchant, A.E. Smith, W.D.M. Rae and J. Zhang, Nucl. Instr. Meth. A351 (1994) 359–370

2.12 1995

1. “Association of the ^{12}C+^{12}C breakup states in ^{24}Mg with the quasi-molecular resonances,” N. Curtis, N.M. Clarke, B.R. Fulton, S.J. Hall, M.J. Leddy, A. St. J. Murphy, J.S. Pople, R.P. Ward, W.N. Catford, G.J. Gyapong, S.M. Singer, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, W.D.M. Rae and P.M. Simmons, Phys. Rev. C51 (1995) 1554–1557

2. “The ^{12}C+^{12}C ‘6–α Chain State’ Resonance,”

S.P.G. Chappell, D.L. Watson, S.P. Fox, C.D. Jones, W.D.M. Rae, P.M. Simmons, M.Freer, B.R. Fulton, N.M. Clarke, N. Curtis, M.J. Leddy, J.S. Pople, S.J. Hall, R.P. Ward, G. Tungate, W.N. Catford, G.J. Gyapong, S.M. Singer and P.H. Regan, Phys. Rev. C51 (1995) 695–700

3. “^{8}Be and α decay of ^{16}O,” M.Freer, N.M. Clarke, N. Curtis, B.R. Fulton, M.J. Leddy, S.J. Hall, J.S. Pople, G. Tungate, R.P. Ward, P.M. Simmons, W.D.M. Rae, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, G.J. Gyapong, S.M. Singer and W.N. Catford, Phys. Rev. C51 (1995) 1682–1692

4. “Evidence for ^{16}O+^{12}C cluster structure in ^{28}Si,”

J.T. Murgatroyd, S.J. Bennett, B.R. Fulton, J.S. Pople, N.S. Jarvis, D.L. Watson, W.D.M. Rae, Y. Chan, D. DiGregorio, J. Scarpaci, J. Suro Perez and R.G. Stokstad, Phys. Rev. C51 (1995) 2230–2232

5. “A schematic model of a 6-α chain state”, W.D.M.Rae and A.C.Merchant, Phys. Rev. Letts. 74 (1995) 4145–4148

6. “Breakup studies with ^{23}Na”, N.S. Jarvis, D.L. Watson, G.J. Gyapong, C.D. Jones, S.J. Bennett, M. Freer, B.R. Fulton, O. Karban, J.T. Murgatroyd, G. Tungate, W.D.M. Rae and A.E. Smith, Phys. Rev. C51 (1995) 2606–2610

7. “Final state interaction in the ^{12}C(^{16}O,^{12}C^{12}C)^{4}He and ^{16}O(^{16}O,^{12}C^{16}O)^{4}He reactions,” M.Freer, N.M. Clarke, R. Le Marechal, G. Tungate, R.P. Ward and W.D.M. Rae, Phys. Rev. C51 (1995) 3174–3183

8. “Relationship between the deformed harmonic oscillator and clustering in light nuclei,” M. Freer, R.R. Betts and A.H. Wuosmaa, Nucl. Phys. A587 (1995) 36–54

9. “Termination of the ^{12}C–^{12}C cluster states revealed in the ^{12}C(^{20}Ne,^{12}C^{12}C)^{8}Be breakup reaction,” J.T. Murgatroyd, S.J. Bennett, N.M. Clarke, B.R. Fulton, G.J. Gyapong, C.D. Jones, D.L. Watson, T.M. Cormier, H. Dejbakhsh, A. Szanto de Toledo and N. Carlin, Nucl. Phys. A587 (1995) 367–

10. “Search for a 7–α chain state,” P.M. Simmons, W.D.M. Rae, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, M.Freer, B.R. Fulton, N.M. Clarke, N. Curtis, M.J. Leddy, J.S. Pople, S.J. Hall, R.P. Ward, G. Tungate, W.N. Catford, G.J. Gyapong, S.M. Singer and P.H. Regan, Phys. Rev. C51 (1995) 3500–3503 11. “Comparison of resonances observed in the ^{12}C(^{24}Mg,^{12}C^{12}C)^{12}C and

12C(^{20}Ne,^{12}C^{12}C)^{8}Be reactions,” M.J. Leddy, S.J. Bennett, N.M. Clarke, M.Freer, B.R. Fulton, J.T. Murgatroyd, G.J. Gyapong, C.D. Jones, D.L. Watson, W.D.M. Rae, A.E. Smith and W.N. Catford, Nucl. Phys. A589 (1995) 363–376

12. “Measurement of low-lying states in ^{9}B,”

M.A. Tiede, K.W. Kemper, N.R. Fletcher, D. Robson, D.D. Caussyn, S.J. Bennett, J.D. Brown, W.N. Catford, C.D. Jones, D.L. Watson and W.D.M. Rae, Phys. Rev. C52 (1995) 1315–1325

13. “Recent Advances in the Study of Nuclear Clusters,”

A. Wuossma, R.R. Betts, M. Freer and B.R. Fulton, Ann. Rev. Nucl. Part. Sci. Astrophys. 45 (1995) 1–41

2.13 1996

1. “Evidence for a highly deformed band in ^{16}O+^{16}O breakup of ^{32}S,” N. Curtis, A. St. J. Murphy, N.M. Clarke, M. Freer, B.R. Fulton, S.J. Hall, M.J. Leddy, J.S. Pople, G. Tungate, R.P. Ward, W.N. Catford, G.J. Gyapong, S.M. Singer, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, W.D.M. Rae, P.M. Simmons, and P.H. Regan, Phys. Rev. C53 (1996) 1804–1810

2. “Search for ^{16}O+ ^{16}O+ ^{16}O structure in ^{48}Cr”, A. St. J. Murphy, R.A. Le Marechal, S.J. Bennett, N.M. Clarke, N. Curtis, M. Freer, B.R. Fulton, S.J. Hall, M.J. Leddy, J.T. Murgatroyd, J.S. Pople, G. Tungate, R.P. Ward, W.D.M. Rae, P.M. Simmons, W.N. Catford, G.J. Gyapong, S.M. Singer, S.P.G. Chappell, S.P. Fox, C.D. Jones, P. Lee and D.L. Watson, Phys. Rev. C53 (1996) 1963–1966

3. “Energy dependence of the ^{16}O( ^{16}O,^{12}C^{∗})^{20}Ne, ^{12}C(^{16}O,^{12}C^{∗})^{16}O and 7Li(^{16}O,^{12}C^{∗})^{11}B reactions”, M. Freer, N.M. Clarke, B.R. Fulton, S.J. Hoad, G.R. Kelly, R.A. Le Marechal, R.P. Ward, C.D. Jones, P. Lee, W.N. Catford and G.J. Gyapong, J. Phys. G22 (1996) 1053–1067

4. “Angular correlations for α–particle decay in the reaction 12C[^{12}C,^{12}C(3_{1}^{−})→^{8}Be(0_{1}^{+}) + α ]^{12}C^{∗} at E_{c.m.} = 32.5 MeV,” S.P.G. Chappell and W.D.M. Rae, Phys. Rev. C53 (1996) 2879–2884

5. “Neutrons from the breakup of ^{19}C,”

F.M. Marqués, E. Liegard, N.A. Orr, J.C. Angélique, L. Axelsson, G. Bizard, W.N. Catford, N.M. Clarke, G. Costa, M. Freer, S. Grévy, D. Guillemand-Mueller, G.J. Gyapong, F. Hanappe, P.G. Hansen, B. Heusch, B. Jonson, C. Le Brun, F.R.

Lecolley, F. Lefebvres, M. Lewitowicz, G. Martinez, A.C. Mueller, T. Nilsson, A. Ninane, G. Nyman, B. Petersen, F. Pougheon, K. Riisager, M.G. Saint-Laurent, Y. Schultz, M. Smedberg, O. Sorlin, L. Stuttgé, D.D. Warner, Phys. Lett. B381 (1996) 407–412

6. “α-cluster structure of the yrast bands of ^{44}Ti,”

J. Zhang, W.D.M. Rae and A.C. Merchant, Phys. Rev. C53 (1996) 515–517

7. “A coupled-channels model for superdeformed prolate nuclei: alpha-chain states in ^{16}O,” A.C. Merchant and W.D.M. Rae, Phys. Rev. C53 (1996) 775–779

8. “Evidence for the 3-fold α-cluster symmetry of ^{12}C,”

W.D.M. Rae, P.E. Fry and A.C. Merchant, J. Phys. G22 (1996) L13–L18.

9. “The analysis of angular correlations in breakup reactions: the effect of coordinate geometries,” M. Freer, Nucl. Instr. Meth. A383 (1996) 463

2.14 1997

1. “Study of alpha-cluster structure using angular momentum projection: Form factors and B(E2) values”, J. Zhang, A.C. Merchant and W.D.M. Rae, Nucl. Phys. A613 (1997) 14–24

2. “Two-dimensional alpha cluster structures in ^{28}Si,”

G.J. Gyapong, N. Curtis, W.N. Catford, S.M. Singer, W.D.M. Rae, M. Freer, S.J. Hall, G.R. Kelly, R.A. Le Marechal, R.P. Ward, S.P.G. Chappell, S.P. Fox, C.D. Jones, P.R. Lee and D.L. Watson, Nucl. Phys. A620 (1997) 55–70

3. “Investigation of the ^{12}C + ^{12}C E_{cm} = 32.5 MeV resonance,”

R.A. Le Marechal, N.M. Clarke, M. Freer, B.R. Fulton, S.J. Hall, S.J. Hoad, G.R. Kelly, R.P. Ward, C.D. Jones, P. Lee and D.L. Watson, Phys. Rev. C55 (1997) 1881–1889

4. “Developments in the study of nuclear clustering in light even-even nuclei”, M. Freer and A.C. Merchant, J. Phys. G23 (1997) 261–323

5. “Calibration of position sensitive detectors with low energy elastic scattering,” S.P.G. Chappell, W.D.M. Rae and P.M. Simmons, Nucl. Instr. Meth. A396 (1997) 383–387 6. “Surface charge flow in two-dimensional position sensitive silicon detectors,”

R.L. Cowin and D.L. Watson, Nucl. Instr. Meth. A399 (1997) 365–381

7. “Search for electromagnetic transitions between ^{12}C–^{12}C cluster states in ^{24}Mg,” F. Haas, A. Elanique, R.M. Freeman, C. Beck, R. Nouicer, D.L. Watson, C. Jones, R. Cowin, P.Lee and Z. Basrak, N. Cimento. 110 (1997) 989

8. “The role of alpha clustering in ^{12}C + ^{12}C scattering,”

W.D.M., Rae, S.P.G. Chappell and P.E. Fry, N. Cimento. A110 (1997) 1001–1006

9. “Cluster states in ^{24}Mg and ^{28}Si,”

B.R. Fulton, M. Freer and J.T. Murgatroyd, N. Cimento. 110 (1997) 1015–1021

10. “Neutron transfer reactions with radioactive beams,” J.S. Winfield, W.N. Catford and N.A. Orr, Nucl. Instr. Meth. A396 (1997) 147

2.15 1998

1. “Can breakup reactions be used to probe high-lying inelastic channels in ^{12}C + 12C scattering?” S.J. Hoad, N.M. Clarke, N. Curtis, M. Freer, B.R. Fulton, S.J. Hall, M.J. Leddy, A.St.J. Murphy, J.S. Pople, G. Tungate, R.P. Ward, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, W.D.M. Rae, P.M. Simmons, W.N. Catford, G.J. Gyapong and S.M. Singer, J. Phys. G24 (1998) 167–173

2. “^{12}C +^{12}C and^{16}O +^{8}Be decay of states excited in the^{12}C(^{16}O,^{24}Mg)α reaction,” M. Freer, N.M. Clarke, B.R. Fulton, J.T. Murgatroyd, A. St.J. Murphy, S.P.G. Chappell, R. Cowin, G. Dillon, D.L. Watson, W.N. Catford, N. Curtis, M. Shawcross and V.F.E. Pucknell. Phys. Rev. C57 (1998) 1277–1289

3. “Search for highly deformed shape isomers in ^{28}Si, ^{24}Mg and ^{20}Ne, linked by successive alpha decay,” G.R Kelly, N.M. Clarke, M. Freer, B.R. Fulton, S.J. Hoad, R.A. Le Marechal and R.P. Ward, Nucl. Phys. A628 (1998) 62–80

4. “Analysis of proton transfer in polarized ^{7}Li scattering by ^{208}Pb at 33 MeV” A.M. Moro, J. Gomez Camacho, I. Martel, M.W. Bailey, K.A. Connell, J.S. Lilley, C.O. Blyth, N.M. Clarke, P.R. Dee, B.R. Fulton, J.A.R. Griffith, S.J. Hall, N. Keeley, K. Rusek, G. Tungate, N.J. Davis, Nucl. Phys. A628 (1998) 203–220 5. “^{16}O + ^{8}Be breakup of ^{24}Mg via the ^{12}C(^{20}Ne,^{16}O ^{8}Be)^{8}Be and ^{12}C(24Mg,^{16}O

8Be)^{12}C reactions,” J.T. Murgatroyd, J.S. Pople, N.M. Clarke, B.R. Fulton, M.J. Leddy, W.N. Catford, S.P. Fox, G.J. Gyapong, C.D. Jones, D.L. Watson, W.D.M. Rae, Y. Chan, R.G. Stokstad and S.J. Bennett , Phys. Rev. C58 (1998) 1569–1575

6. “High spin resonances in ^{12}C + ^{12}C scattering,”

S.P.G. Chappell, W.D.M. Rae, C.A. Bremner, G.K. Dillon, D.L. Watson, B. Greenhalgh, R.L. Cowin, M. Freer and S.M. Singer, Phys. Lett. B444 (1998) 260

2.16 1999

1. “Excitation function measurements for the ^{12}C(^{20}Ne,^{12}C[0_{2}^{+}])^{20}Ne reaction,” S.J. Hoad, N.M. Clarke, M. Freer, B.R. Fulton, P. Fry, W.D.M. Rae and A.H. Wuosmaa, Phys. Rev. C59 (1999) 1456

2. “Structures in the ^{12}C(^{16}O,^{12}C(0_{2}^{+}) → 3α)^{16}O reaction in the region E_{cm} = 15 to 34 MeV, S.P. Fox, D.L. Watson, C.D. Jones, S.P. Chappell, P. Lee, W.D.M. Rae, M. Freer, S.J.A. Hall, G. Kelly, R. Le Marechal, R.P. Ward, W.N. Catford, G.J. Gyapong and S. Singer, Phys. Rev. C60 (1999) 064620

3. “Core excitation in ^{11}Be_{gs} via the p(^{11}Be,^{10}Be)d reaction,”

S. Fortier, S. Pita, J.S. Winfield, W.N. Catford, N.A. Orr, J. Van de Wiele, Y. Blumenfeld, R. Chapman, S.P.G. Chappell, N.M. Clarke, N. Curtis, M. Freer, S. Galès, K.L. Jones, H. Langevin-Joliot, H. Laurent, I. Lhenry, J.M. Maison, P. Roussel-Chomaz, M. Shawcross, M. Smith, K. Spohr, T. Suomijarvi and A. de Vismes, Phys. Lett. B461 (1999) 22–27

4. “Exotic molecular states in ^{12}Be,”

M. Freer, J.C. Angélique, L. Axelsson, B. Benoit, U. Bergmann, W.N. Catford, S.P.G. Chappell, N.M. Clarke, N. Curtis, A. D’Arrigo, E. de Goes Brennard, O. Dorvaux, B.R. Fulton, G. Giardina, C. Gregori, S. Grévy, F. Hanappe, G. Kelly, M. Labiche, C. Le Brun, S. Leenhardt, M. Lewitowicz, K. Markenroth, F.M. Marqués, M. Motta, J.T. Murgatroyd, T. Nilsson, A. Ninane, N.A. Orr, I. Piqueras, M.G. Saint Laurent, S.M. Singer, O. Sorlin, L. Stuttgé and D.L. Watson, Phys. Rev. Lett. 82 (1999) 1383–1386

5. “Structure of ^{11}Be from the (d,p) reaction in inverse kinematics,”

J.S. Winfield, S. Fortier, W.N. Catford, S. Pita, N.A. Orr, Y. Blumenfeld, R. Chapman, S.P.G. Chappell, N.M. Clarke, N. Curtis, M. Freer, S. Galès, K.L. Jones, H. Langevin-Joliot, H. Laurent, I. Lhenry, J.-M. Maison, P. Roussel-Chomaz, M.

Shawcross, M. Smith, K. Spohr, T. Suomijarvi and A. de Vismes, J. Phys. G25 (1999) 755–757

6. “Fission dynamics for capture reactions in ^{58,64}Ni + ^{208}Pb systems: New results in terms of thermal energy and neutron multiplicity correlated distributions,” L. Donadille, E. Liatard, B. Benoit, F. Hanappe, L. Stuttge, G. Rudolf, E.M. Kozulin, Yu.A. Lazarev, P. Desesquelles, L.A. Litnevsky, B. Bilwes, J.-F. Bruandet, G.J. Costa, O. Dorvaux, F. Farget, J. Fayot, G. Guillaume, A. Huck, I. Itkis, M.G. Itkis, S. Jokic, N.A. Kondratev, F.-R. Lecolley, J.-F. Lecolley, Yu.Ts. Oganessian, G. Perrin, D. Santos, F. Schussler, S. Tomasevic, J.-B. Viano and B. Vignon, Nucl. Phys. A656 (1999) 259

7. “A new detector array for charged particle spectroscopy,”

R.L. Cowin, D.L. Watson, S.P.G. Chappell, N.M. Clarke, M. Freer, B.R. Fulton, R.A. Cunningham, N. Curtis, G.K. Dillon, J. Lilley, C.D. Jones, P. Lee and W.D.M. Rae, Nucl. Instr. Meth. A423 (1999) 75–91

8. “Deviation from tidal symmetry for scattering of polarized^{23}Na from the deformed nucleus ^{27}Al,” N.J. Davis, H.D. Choi, R.P. Ward, P.J. Woods, D.G. Ireland, K. Livingston, R.D. Page, P.J. Sellin, A.C. Shotter, J.A.R. Griffith, S.J. Hall, P.R. Dee, I. Martel-Bravo, J.M. Nelson and K. Rusek, Phys. Rev. C60 (1999) 034602

9. “Clustering in nuclei: nuclear chains, nuclear molecules and other exotic states of nuclei,” B.R. Fulton, Contemp. Phys. 40 (1999) 299–311

2.17 2000

1. “Mechanism of the^{12}C(^{24}Mg,^{12}C^{12}C)^{12}C breakup reaction: direct or compound?,” S.M.Singer, W.N. Catford, G.J. Gyapong, M. Shawcross, N.M. Clarke, N. Curtis, M. Freer, B.R. Fulton, S.J. Hall, R.A. Le Marechal, M.J. Leddy, J.S. Pople, G. Tungate, R.P Ward, W.D.M. Rae, P.M. Simmons, S.P.G. Chappell, S.P. Fox, C.D. Jones, P. Lee and D.L. Watson, Phys. Rev. C62 (2000) 054609

2. “Two-neutron interferometry as a probe of the nuclear halo,”

F.M. Marqués, M. Labiche, N.A. Orr, J.C. Angélique, L. Axelsson, B. Benoit, U.C. Bergmann, M.J.G. Borge, W.N. Catford, S.P.G. Chappell, N.M. Clarke, G. Costa, N. Curtis, A. D’Arrigo, F. de Oliveira Santos, E. de Goes Brennard, O. Dorvaux, M. Freer, B.R. Fulton, G. Giardina, C. Gregori, S. Grévy, D. Guillemaud-Mueller, F. Hanappe, B. Heusch, B. Jonson, C. Le Brun, S. Leenhardt, M. Lewitowicz, M.J. Loṕez, K. Markenroth, M. Motta, A.C. Mueller, T. Nilsson, A. Ninane, G.

Nyman, I. Piqueras, K. Riisager, M.G. Saint Laurent, F. Sarazin, S.M. Singer, O. Sorlin and L. Stuttgé, Phys. Lett. B476 (2000) 219–225

3. “Target dependence in the resonant ^{12}C + ^{12}C breakup of excited states in ^{24}Mg,” N. Curtis, M. Shawcross, W.N. Catford, B.R. Fulton, N.M. Clarke, S.J. Hall, J.T. Murgatroyd, S.P.G. Chappell, R. Cowin, G. Dillon and D.L. Watson, Phys. Rev. C61 (2000) 064606

4. “Reaction mechanism populating ^{12}C + ^{16}O breakup states in ^{28}Si,”

N. Curtis, M. Shawcross, W.N. Catford, B.R. Fulton, N.M. Clarke, S.J. Hall, J.T. Murgatroyd, S.P.G. Chappell, R. Cowin, G. Dillon and D.L. Watson, Phys. Rev. C62 (2000) 034603

5. “One-neutron removal reactions on neutron-rich psd-shell nuclei,”

E. Sauvan, F. Carstoiu, N.A. Orr, J.C. Angélique, W.N. Catford, N.M. Clarke, M. Mac Cormick, N. Curtis, M. Freer, S. Grévy, C. Le Brun, E. Liégard, M. Lewitowicz, F.M. Marqués, G. Martinez, P. Roussel-Chomaz, M.G. Saint-Laurent, M. Shawcross and J.S. Winfield, Phys. Lett. B491 (2000) 1

6. “Detection of particle unbound states using a single Gas-Si-CsI hybrid telescope,” M. Shawcross, W.N. Catford, N. Curtis, N.M. Clarke, S.P.G. Chappell, B.R. Fulton and D.L. Watson, Nucl. Instr. Meth. A450 (2000) 373

7. “Elastic scattering of ^{16}O + ^{16}O at energies E/A between 5 and 8 MeV,” M.P. Nicoli, F. Haas, R.M. Freeman, N. Aissaoui, C. Beck, A. Elanique, R. Nouicer, A. Morsad, S. Szilner, Z. Basrak, M.E. Brandan and G.R. Satchler, Phys. Rev. C60 (2000) 064608

8. “Detailed study and mean field interpretation of ^{16}O + ^{12}C elastic scattering at seven medium energies,” M.P. Nicoli, F. Haas, R.M. Freeman, S. Szilner, Z. Basrak, A. Morsad, G.R. Satchler and M.E. Brandan, Phys. Rev. C61 (2000) 034609

9. “Data acquistion system for MEGHA,”

S.P.G. Chappell, R.A. Hunt, D. Smith, W.D.M Rae, N.M. Clarke, M. Freer, B.R. Fulton, S.S. Jagpal, S.M. Singer and D.L. Watson, Nucl. Instr. Meth. A450 (2000) 399–414

2.18 2001

1. “Search for molecular states in ^{16}C,”

P.J. Leask, L. Achouri, J.C. Angélique, B. Benoit, C. Bremner, W.N. Catford, R. Chapman, S.P.G. Chappell, N.M. Clarke, E. de Góes Brennard, L. Donadille, M.

Freer, B.R. Fulton, S. Grévy, F. Hanappe, K.L. Jones, J.L. Lecouey, R. Lemmon, M. Lewitowicz, X. Liang, F.M. Marqués, M.P. Nicoli, A. Ninane, N.A. Orr, M. Shawcross, S.M. Singer, M. Smith, K. Spohr and L. Stuttgé, J. Phys. G27 (2001) B9–B14

2. “The ^{12}C(^{16}O,^{12}C ^{12}C)α reaction at E_{beam} = 75 MeV,”

M. Freer, M.P. Nicoli, N.M. Clarke, B.R. Fulton, G.R. Kelly, J.T. Murgatroyd and S.M. Singer, Nucl. Phys. A694 (2001) 147–156

3. “A new coupling potential for the analysis of deformed light heavy-ion reactions,” I. Boztosun and W.D.M. Rae, Phys. Lett. B518 (2001) 229

4. “Analysis of the ^{12}C + ^{12}C reaction using a new type of coupling potential,” I. Boztosun and W.D.M. Rae, Phys. Rev. C63 (2001) 054607

5. “α breakup of ^{6}Li and ^{7}Li near the coulomb barrier,”

G.R. Kelly, N.J. Davis, R.P. Ward, B.R. Fulton, G. Tungate, N. Keeley, K. Rusek, E.E. Bartosz, P.D. Cathers, D.D. Caussyn, T.L. Drummer and K.W. Kemper, Phys. Rev. C63 (2001) 024601

6. “Halo structure of ^{14}Be,”

M. Labiche, F.M. Marqués, N.A. Orr, J.C. Angélique, L. Axelsson, B. Benoit, U. Bergmann, M.J.G. Borge, W.N. Catford, S.P.G. Chappell, N.M. Clarke, G. Costa, N. Curtis, A. D’Arrigo, E. de Goes Brennard, O. Dorvaux, M. Freer, B.R. Fulton, G. Giardina, C. Gregori, S. Grévy, D. Guillemaud-Mueller, F. Hanappe, B. Heusch, K.L. Jones, B. Jonson, C. Le Brun, S. Leenhardt, M. Lewitowicz, M.J. Lopez, K. Markenroth, M. Motta, A.C. Mueller, T. Nilsson, A. Ninane, G. Nyman, F. de Oliveira, I. Piqueras, K. Riisager, M.G. Saint Laurent, F. Sarazin, S.M. Singer, O. Sorlin, L.Stuttgé, Phys. Rev. Lett. 86 (2001) 600

7. “Helium breakup states in ^{10}Be and ^{12}Be,”

M. Freer, J.C. Angélique, L. Axelsson, B. Benoit, U. Bergmann, W.N. Catford, S.P.G. Chappell, N.M. Clarke, N. Curtis, A. D’Arrigo, E. de Goes Brennard, O. Dorvaux, B.R. Fulton, G. Giardina, C. Gregori, S. Grévy, F. Hanappe, G. Kelly, M. Labiche, C. Le Brun, S. Leenhardt, M. Lewitowicz, K. Markenroth, F.M. Marqués, M. Motta, J.T. Murgatroyd, T. Nilsson, A. Ninane, N.A. Orr, I. Piqueras, M.G. Saint Laurent, S.M. Singer, O. Sorlin, L. Stuttgé and D.L. Watson, Phys. Rev. C63 (2001) 034301 Phys. Rev. C64 (2001) 019904 (Erratum)

8. “Breakup measurements of particle unbound states in ^{10}B,”

P.J. Leask, M. Freer, N.M. Clarke, B.R. Fulton, C.D. Freeman, S.M. Singer, W.N. Catford, N. Curtis, K.L. Jones, R.L. Cowin, D.L. Watson, R.P. Ward, N.A. Orr and V.F.E. Pucknell, Phys. Rev. C63 (2001) 034307

9. “Angular correlations from the ^{12}C + ^{16}O breakup of ^{28}Si and ^{12}C + ^{12}C breakup of ^{24}Mg,” M. Shawcross, N. Curtis, W.N. Catford, N.M. Clarke, B.R. Fulton, S.J. Hall, J.T. Murgatroyd, S.P.G. Chappell, R. Cowin, G. Dillon and D.L. Watson, Phys. Rev. C63 (2001) 034311

10. “Evidence for a highly deformed band in ^{12}C + ^{16}O breakup of ^{28}Si,” N.I. Ashwood, J.T. Murgatroyd, N.M. Clarke, M. Freer, B.R. Fulton, A. St.J. Murphy, S.P.G. Chappell, R.L. Cowin, G.K. Dillon, D.L. Watson, W.N. Catford, N. Curtis, M. Shawcross and V. Pucknell, Phys. Rev. C63 (2001) 034315

11. “The ^{12}C + ^{12}C and ^{16}O + ^{8}Be decay of ^{24}Mg states populated in the ^{12}C(^{20}Ne, 24Mg^{∗})^{8}Be reaction,” M. Freer, J.T. Murgatroyd, S.M. Singer, N. Curtis, D.J. Henderson, D.J. Hofman and A.H. Wuosmaa, Phys. Rev. C63 (2001) 034317

12. “Single-neutron transfer from ^{11}Be_{gs} via the (p,d) reaction with a radioactive beam,” J.S. Winfield, S. Fortier, W.N. Catford, S. Pita, N.A. Orr, J. Van de Wiele, Y. Blumenfeld, R. Chapman, S.P.G. Chappell, N.M. Clarke, N. Curtis, M. Freer, S. Galès, H. Langevin-Joliot, H. Laurent, I. Lhenry, J.M. Maison, P. Roussel-Chomaz, M. Shawcross, K. Spohr, T. Suomijarvi and A. de Vismes, Nucl. Phys. A683 (2001) 48–78

13. “Decay angular correlations and spectroscopy for ^{10}Be^{∗} → ^{4}He + ^{6}He,” N. Curtis, D.D. Caussyn, N.R. Fletcher, F. Maréchal, N.Fay and D. Robson, Phys. Rev. C64 (2001) 044604

14. “Nuclear Molecules,”

M. Freer, Nucl. Phys. A685 (2001) 146c

15. “The ‘How and Why’ of nucleon transfer reactions with radioactive beams,” W.N. Catford, Acta. Phys. Pol. B32 (2001) 1049

16. “A determination of the ^{6}He + p interaction potential,”

A.de. Vismes, P. Roussel-Chomaz, W. Mittig, A. Pakou, N. Alamanos, F. Auger, J.-C. Angelique, J. Barrette, A.V. Belozyorov, C. Borcea, W.N. Catford, M.D. Cortina-Gil, Z. Dlouhy, A. Gillibert, V. Lapoux, A. Lepine-Szily, S.M. Lukyanov, F. Marie, A. Musumarra, F. de Oliveira, N.A. Orr, S. Ottini-Hustache, Y.E. Penionzhkevich, F. Sarazin, H. Savajols and N. Skobelev, Phys. Lett. B505 (2001) 15 17. “Radiative proton capture on ^{6}He,”

E. Sauvan, F.M. Marques, H.W. Wilschut, N.A. Orr, J.C. Angelique, C. Borcea, W.N. Catford, N.M. Clarke, P. Descouvemont, J. Diaz, S. Grevy, A. Kugler, V. Kravchuk, M. Labiche, C. Le Brun, E. Lienard, H. Lohner, W. Mittig, R.W. Ostendorf, S. Pietri, P. Roussel-Chomaz, M.G. Saint laurent, H. Savajols, V. Wagner and N. Yahlali, Phys. Rev. Lett. 87 (2001) 042501

3 Papers Accepted for Publication

1. “New findings for the ^{12}C + ^{12}C reaction,”

I. Boztosun and W.D.M. Rae, Accepted for publication in Physics of Atomic Nuclei

2. “A global potential analysis of the^{16}O +^{28}Si reaction using a new yype of coupling potential,” I. Boztosun and W.D.M. Rae, Accepted for publication in Phys. Rev. C.

3. “Three-body correlations in Borromean haloes,”

F.M. Marqués, M. Labiche, N.A. Orr, J.C. Angélique, L. Axelsson, B. Benoit, U.C. Bergmann, M.J.G. Borge, W.N. Catford, S.P.G. Chappell, N.M. Clarke, G. Costa, N. Curtis, A. D’Arrigo, F. de Oliveira Santos, E. de Goes Brennard, O. Dorvaux, M. Freer, B.R. Fulton, G. Giardina, S. Grévy, D. Guillemaud-Mueller, F. Hanappe, B. Heusch, B. Jonson, C. Le Brun, S. Leenhardt, M. Lewitowicz, M.J. Lopez, K. Markenroth, M. Motta, A.C. Mueller, T. Nilsson, A. Ninane, G. Nyman, I. Piqueras, K. Riisager, M.G. Saint Laurent, F. Sarazin, S.M. Singer, O. Sorlin, L.Stuttgé, Accepted for publication in Phys. Rev. C.

4. “Systematics of breakup yield for weakly bound nuclei,”

B.R. Fulton, R.J. Woolliscroft, N.M. Clarke and M. Freer, Accepted for publication in Phys. Rev. C.

4 Papers Recently Submitted

1. “Near threshold three-body final states in ^{7}Li + ^{7}Li reactions at E_{lab} = 34 MeV,” J.A. Liendo, N. Curtis, D.D. Caussyn, N.R. Fletcher and T. Kurtukian-Nieto, Submitted to Phys. Rev. C.

5 Papers in Preparation

1. “Observation of an octupole rotational band populated in the ^{14}C + α decay of 18O and evidence for the ^{18}O + α decay of ^{22}Ne,” N. Curtis, D.D. Caussyn, C. Chandler, M.W. Cooper, N.R. Fletcher, R.W. Laird and J. Pavan, To be submitted to Phys. Rev. C.

2. ”Search for analogue molecular states in ^{16}C,”

B. Greenhalgh, B.R. Fulton, D.L. Watson, N.M. Clarke, L. Donadille, M. Freer, P.J. Leask, W.N. Catford, K.L. Jones, and D. Mahboub

6 Invited Talks at Conferences

1. “The Study of Cluster States in Light Nuclei via Heavy-ion-alpha Correlation Measurements,” W.D.M. Rae, Proc. 4th Int. Conf. on Clustering Aspects of Nuclear Structure and Nuclear Reactions, Chester, 1984 (D. Reidel, publisher).

2. “Clustering aspects of nuclear structure and reactions,”

W.D.M. Rae, Int. Nucl. Phys. Conf., Harrogate, 1986, Inst. of Phys. Conf. Series, 86, p. 189–203

3. “Symmetric Fission of sd Shell Nuclei,”

American Physical Society Spring Meeting, Crystal City VA (USA), 1987.

4. “Clustering Phenomena and Shell Effects in Nuclear Structure and Reactions,” W.D.M. Rae, Proc. of the X^{th} Symposium on Nuclear Physics, Oaxtepec, Mexico, Notas de Fisica 10 (1987) 283

5. “Continuum Breakup Spectroscopy in sd Shell Nuclei,”

W.D.M. Rae, Nuclear Breakup and Recoil Spectroscopy Study Weekend, Daresbury (UK), October 1987.

6. “Clustering and Shape Information from Breakup Reactions,”

IOP Nuclear Physics Conference, Manchester (UK), March 1988.

7. “Relationships between Clustering and Shell Structure in Nuclei and Nuclear Reactions,” W.D.M. Rae, Fifth International Conference on Clustering Aspects in Nuclear and Subnuclear Systems, Kyoto (Japan), July 1988.

8. “^{16}O+ ^{8}Be Cluster Structure in ^{24}Mg,”

B.R. Fulton, Fifth International Conference on Clustering Aspects in Nuclear and Subnuclear Systems, Kyoto (Japan), July 1988.

9. “The breakup of sd-shell nuclei,”

International symposium on the developments in nuclear cluster dynamics, Sapporo, Japan, July 1988.

10. “Spins and Structure of ^{12}C — ^{12}C Resonances,” Proc. of the JAERI Int. Symp. on Heavy Ion Reaction Dynamics in the Tandem Energy Region, Hitachi, Japan (1988) 197 11. “Clustering in Nuclei,”

W.D.M. Rae, Proc. International Nuclear Physics Conference, Sao Paulo, Brazil, 1989, (World Scientific), Vol. 2, p. 299

12. “Relationships between clustering and shell structure in nuclei and nuclear reactions”, W.D.M. Rae, Proc. Fifth Int. Conf. on Clustering Aspects in Nuclear and Subnuclear Physics, Kyoto, Japan, 1988, J. Phys. Soc. Jpn. 58 (1989) Suppl. p.77–87

13. “^{16}O + ^{8}Be Cluster Structure in ^{24}Mg,”

B.R. Fulton, Proc. Fifth Int. Conf. on Clustering Aspects in Nuclear and Subnuclear Physics, Kyoto, Japan, 1988, J. Phys. Soc. Jpn. 58 (1989) Suppl. p.590.

14. “Spin measurement for symmetric fissioning states in ^{24}Mg,”

M. Freer, Interaction of nuclear structure and heavy ion reaction dynamics, Institute of Physics Conference Series, 109 (1990) 243

15. “Experimental techniques for ^{1}H(^{23}Na,^{12}C)^{12}C using thick targets,”

N.M. Clarke, Workshop on Resonant Scattering, Brussels, 9 December 1992.

16. “Studies of clustering in light nuclei, European forum on nuclear reaction mechanism studies, Legnaro, Italy, December 1992.

17. “Clustering in Nuclei: the Charissa Story,”

Requiem for an Accelerator, Daresbury, March 1993.

18. “Nuclear cluster states”, in Nuclear and Particle Physics 1993, ed. I.J.D. MacGregor and A.T. Doyle, Institute of Physics Conference Series,133 (1993) 35–49

19. “Recent studies of clustering in nuclei,”

Workshop on open problems in heavy ion reaction dynamics at Vivitron energies, Strasbourg, 6 May 1993.

20. “CHARISSA: The UK charged particle array,”

D.L. Watson, Workshop on open problems in heavy ion reaction dynamics at Vivitron energies, Strasbourg, 6 May 1993.

21. “Nuclear Molecules and Nuclear Clusters,”

Second International Conference on atomic and nuclear clusters, Santorini, Greece, 28 June–2 July 1993.

22. “Alpha chain states in 4N-nuclei”, Second International Conference on atomic and nuclear clusters, Santorini, Greece, 28 June–2 July 1993.

23. “Chain states in light nuclei”, M. Freer, Gordon Research Conference, New London, New hampshire, July 1993.

24. “Correlations between alpha particles: chain like nuclei”, Ninth Amsterdam Mini-Conference, NIKHEF, 9–10 November 1993.

25. “Study of clustering through breakup reactions”, W.D.M. Rae, invited talk at the Institute of Physics Meeting on Nuclear and Atomic/Molecular Clustering, Oxford, 9 March 1994.

26. “Clustering in Light Nuclei,”

M. Freer, invited talk at the Institute of Physics Annual Congress, Brighton, U.K. 11–14 April 1994.

27. “Nuclear Clusters,”

B.R. Fulton, invited talk at BrisPhys ’94, Sixth AAPC and Eleventh Australian Institute of Physics Congress, Brisbane, 4–8 July 1994.

28. “Probing cluster structures through breakup reactions,”

M. Freer, invited talk at the “6th Int. conf. on clusters in nuclear structure and dynamics”, Strasbourg, France, 6–9 September 1994.

29. “Superdeformed α Cluster States and Heavy Ion Resonances,”

W.D.M. Rae, invited talk at the “6th Int. conf. on clusters in nuclear structure and dynamics”, Strasbourg, France, 6–9 September 1994.

30. “Clusters in nuclear structure and dynamics: the Experimental Point of View,” B.R. Fulton, invited talk at the “6th Int. conf. on clusters in nuclear structure and dynamics”, Strasbourg, France, 6–9 September 1994.

31. “Alpha and exotic cluster structure in heavy nuclei, A.C. Merchant, invited talk at the EPSRC Nuclear Physics Study Weekend, Abingdon, October 1995

32. “Experimental evidence of nuclear cluster states,”

W.N. Catford, invited talk at the EPSRC Nuclear Physics Study Weekend, Abingdon, October 1995

33. “Cluster measurements with the Charissa array,” B.R. Fulton, invited talk at the Joint Japan-US seminar on clustering in nuclear and mesoscopic systems, Hawaii, 14–16 December 1995.

34. “The 6-alpha chain state; does it exist?,”

B.R. Fulton, invited talk at the International conference on nuclear dynamics at long and short distances, Angra dos Reis, Brazil, 8–12 April 1996.

35. “Breakup studies with Charissa,”

N. Curtis, invited talk at the Institute of Physics Half Day Meeting, Glasgow, November 1996.

36. “The shape of the shape eigenstate,”

W.D.M. Rae, invited talk at the Int. Symp. on Nucl. Phys., Bombay, December 1995, Department of atomic energy proceedings p. 94.

37. “Clustering in sd shell nuclei,”

M. Freer, invited talk at the EPSRC Nuclear Physics Study Weekend, Birmingham, October 1996

38. “The role of α-clustering in ^{12}C + ^{12}C scattering,” S.P.G. Chappell, invited talk at the Institute of Physic Nuclear Physics Conference, York 7 – 9th April 1997.

39. “Cluster states in ^{24}Mg and ^{28}Si,”

B.R. Fulton, invited talk at the International Workshop on New Ideas on Clustering in Nuclear and Atomic Systems, 9-13 June 1997, Rauischlozhausen Castle, Germany

40. “The role of α-clustering in ^{12}C + ^{12}C scattering,” W.D.M. Rae, invited talk at the International Workshop on New Ideas on Clustering in Nuclear and Atomic Systems, 9-13 June 1997, Rauischlozhausen Castle, Germany

41. “Resonances or Fluctuations? The challenge,” W.D.M. Rae, invited talk at the EPSRC Nuclear Physics Study Weekend, Birmingham, October 1997 42. “Recent results from the Charissa collaboration,”

B.R. Fulton, invited talk at the Institute of Physics Nuclear Physics Conference, Salford, April 1999

43. “Charged particle spectroscopy at GANIL,”

M. Freer, invited talk at the Institute of Physics Half Day Meeting, Edinburgh, October 1998.

44. “The six-alpha chain state story,”

B.R. Fulton,” inivted talk at the 7th International Conference on clusters in Nuclear Structure and Nuclear Dynamics, Rab, Croatia, June 1999

45. “The development and applications of scintillator-photodiode detectors,” N.M. Clarke, invited talk and proc. of “The nucleus: new physics for the new millenium,” Cape Town, S. Africa 18-22 January 1999, Ed. F.D. Smit, R. Lindsay and S.V. Fortsch, Kluwer Academic/Plenum Publishers 2000

46. “Clustering from N/Z = 1 to N/Z = 2,”

W.N. Catford, invited talk and proc. of “The nucleus: new physics for the new millenium,” Cape Town, S. Africa 18-22 January 1999, Ed. F.D. Smit, R. Lindsay and S.V. Fortsch, Kluwer Academic/Plenum Publishers 2000

47. “SIRIUS,”

M. Freer, invited talk at the Beta decay workshop, Strasbourg, France, March 1999

48. “Recent results from the Charissa collaboration,”

B.R. Fulton, invited talk at the Institute of Physics nuclear and particle physics conference, Salford, April 1999

49. “By transfer, to the haloes,”

W.N. Catford, invited talk at the Institute of Physics nuclear and particle physics conference, Salford, April 1999

50. “Clustering in exotic nuclei,”

M. Freer, invited talk at the GANIL Colloque, Biarritz, May 1999

51. “^{6}He + ^{6}He breakup states in ^{12}Be,”

M. Freer, invited talk and Proc. 7th Intern. Conf. on Clustering Aspects of Nuclear Structure and Dynamics, Rab, Island of Rab, Croatia, 14–19 June 1999, Ed. M. Korolija, Z. Basrak, R. Caplar, World Scientific (2000)

52. “The six alpha chain state story,”

B.R. Fulton, invited talk and Proc. 7th Intern. Conf. on Clustering Aspects of Nuclear Structure and Dynamics, Rab, Island of Rab, Croatia, 14–19 June 1999, Ed. M. Korolija, Z. Basrak, R. Caplar, World Scientific (2000)

53. “Nuclear rainbows,”

M.-P. Nicoli, invited talk and Proc. 7th Intern. Conf. on Clustering Aspects of Nuclear Structure and Dynamics, Rab, Island of Rab, Croatia, 14–19 June 1999, Ed. M. Korolija, Z. Basrak, R. Caplar, World Scientific (2000)

54. “Coupled-channels calculations for the scattering of deformed light nuclei,” I. Boztosun and W.D.M. Rae, invited talk and Proc. 7th Intern. Conf. on Clustering Aspects of Nuclear Structure and Dynamics, Rab, Island of Rab, Croatia, 14–19 June 1999, Ed. M. Korolija, Z. Basrak, R. Caplar, World Scientific (2000)

55. “Experimental studies of exotic nuclei at GANIL,”

N.M. Clarke, invited talk, IOP Nuclear Physics Conference, Birmingham 2000

56. “Nuclear molecules,”

M. Freer, invited talk at Nucleus - Nucleus 2000, Strasbourg, France 2000 and Nucl. Phys. A685 (2001) 146c

57. “Clustering in nuclei: nuclear chains, nuclear molecules and other exotic states of nuclei,” B.R. Fulton, invited talk at the TRIUMF Laboratory, October 2000

58. “Nucleosynthesis in supernovae - what can we learn at ISAC-II?,”

B.R. Fulton, invited talk at the TRIUMF meeting on opportunities at ISAC-II, 2000

59. “The experimental study of exotic nuclear states using multi-detector arrays,” D.L. Watson, invited talk at the International workshop on physics with multi-detector arrays, Calcutta, India, November 2000

60. “Nuclear molecules,”

M. Freer, invited talk at the RIKEN symposium on Physics at the drip lines, February 2001 and RIKEN Review 39 (2001) 109 61. “Overview to clustering,”

W.N. Catford, invited talk at “Haloes, Peaux et Driplines,”, GANIL, France, May 2001.

62. “Clustering and molecular states in exotic nuclei,” M. Freer, invited talk at “Haloes, Peaux et Driplines,”, GANIL, France, May 2001.

63. “Recent Results on ^{10,12}Be,”

N. Curtis, invited talk at “Haloes, Peaux et Driplines,”, GANIL, France, May 2001.

64. “From clusters to molecules,”

M. Freer, invited talk at Future of nuclear structure and gamma spectroscopy with stable beams, Strasbourg, France, June 2001

65. “Reactions and breakup with ^{9}Be, B.R. Fulton, invited talk at the International workshop on the physics of halo nuclei, Trento, October 2001

7 Conference Contributions

1. “A Study of^{16}O and^{20}Ne using the techniques of Resonant Particle Spectroscopy,” W.D.M. Rae, S.C. Allcock, A.E. Smith and B.R. Fulton, Proc. of 1984 INS-RIKEN Int. Symp. on Heavy Ion Physics.

2. “A Study of^{16}O and^{20}Ne using the techniques of Resonant Particle Spectroscopy,” W.D.M. Rae, S.C. Allcock, A.E. Smith and B.R. Fulton, IOP Nuclear Physics Conference, Bradford, 1984.

3. “A Study of^{16}O and^{20}Ne using the techniques of Resonant Particle Spectroscopy,” W.D.M. Rae, S.C. Allcock, A.E. Smith and B.R. Fulton, 4th Int. Conf. on Clustering Aspects of Nuclear Structure and Nuclear Reactions, Chester, 1984.

4. “A Search for Heavy Ion resonances at A=32 using ^{12}C + ^{20}Ne,”

N.R. Fletcher, D.L. Gay, W.D.M. Rae, A.E. Smith and S.C. Allcock, Bull. Am. Phys Soc. 36 (1985) 1768

5. “High Spin α-cluster Resonances in ^{16}O,”

W.D.M. Rae, S.C. Allcock, S. Marsh and B.R. Fulton, Proc. of the 4th Int. Conf. on Nuclear Reaction Mechanisms, Varenna, 1985, pp. 517-528

6. “Symmetric Fission of ^{24}Mg,”

B.R. Fulton, S.J. Bennett, C.A. Ogilvie, J.S. Lilley, D.W. Banes, W.D.M. Rae, S.C. Allcock, R.R. Betts and A.E. Smith, Proc. of the ANL symposium on the Many Facets of Heavy Ion Fusion Reactions, Argonne, March 1986.

7. “Symmetric Fission of ^{24}Mg,”

B.R. Fulton, S.J. Bennett, C.A. Ogilvie, J.S. Lilley, D.W. Banes, W.D.M. Rae, S.C. Allcock, R.R. Betts and A.E. Smith, Proc. of the Int. Nucl. Phys. Conf., Harrogate, 1986.

8. “Spin Information for states in ^{24}Mg at E_{x} > 12 MeV via (α, α) Correlation Measurements,” A.E. Smith, S.C. Allcock, P.R. Keeling and W.D.M. Rae, Proc. of the Int. Nucl. Phys. Conf., Harrogate, 1986.

9. “The Cranked Cluster Model and its Application to ^{24}Mg”, S. Marsh and W.D.M. Rae, Proc. of the Int. Nucl. Phys. Conf., Harrogate, 1986.

10. “High Spin Alpha-Cluster Resonances in ^{16}O and ^{20}Ne,” S.C. Allcock, W.D.M. Rae, P.R. Keeling, S. Marsh, A.E. Smith, B.R. Fulton and D.W. Banes, Proc. of the Int. Nucl. Phys. Conf., Harrogate, 1986.

11. “A Study of ^{12}C–^{12}C Resonances via the ^{12}C(^{12}C,αα)^{16}O Reaction at E_{CM} = 19.32 MeV,” W.D.M. Rae, P.R. Keeling and S.C. Allcock, Proc. of the Int. Nucl. Phys. Conf., Harrogate, 1986.

12. “Spin and Structure of the Broad ^{12}C — ^{12}C Resonance at E_{CM} = 25 MeV via the ^{12}C,α^{8}Be)^{12}C Reaction,” P.R. Keeling, W.D.M. Rae and A.E. Smith, Proc. of the I.O.P. Conference, 6–8 April 1987, University of Birmingham.

13. “Augmenting Camac with Links and Transputers,”

J.D. Brown, R.A. Hunt, W.D.M. Rae, S.J. Bennett, M. Freer and B.R. Fulton, International Symposium on Nuclear Electronics and Seminar CAMAC-90, 25–28 September 1990, Warsaw.

14. “Low Lying Excited States in ^{9}B”, M.A. Tiede, D.D. Caussyn, N.R. Fletcher, J.K. Harris, K.W. Kemper, S.J. Bennett, C. Jones, D.L. Watson, W.D.M. Rae, J.D. Brown and W.N. Catford, American Physical Society (Dec 1991)

15. “Study of ^{12}C+^{12}C breakup of ^{24}Mg following alpha–transfer onto ^{20}Ne”, J.T. Murgatroyd, S.J. Bennett, B.R. Fulton, W.D.M. Rae, G.J. Gyapong, N.S. Jarvis, C.D. Jones, D.L. Watson, Y. Chan, D. DiGregorio, J. Scarpaci, J. Suro Perez, R.G. Stokstad, T.M. Cormier, H. Dejbakhsh, N. Carlin and A. Szanto de Toledo, Scientific Program of the North-West Europe Nuclear Physics Conference, Edinburgh (1992), p.44.

16. “^{16}O+^{8}Be breakup of ^{24}Mg”, J.S. Pople, B.R. Fulton, S.J. Bennett, J.T. Murgatroyd, M.J. Leddy, D.L. Watson, G.J. Gyapong, C.D. Jones, T. Davinson and W.D.M. Rae, Scientific Program of the North-West Europe Nuclear Physics Conference, Edinburgh (1992), p.45.

17. “Further Investigations into Breakup States in ^{24}Mg”, M.J. Leddy, B.R. Fulton, S.J. Bennett, N.M. Clarke, M. Freer, J.T. Murgatroyd, G.J. Gyapong, C.D. Jones, D.L. Watson, W.D.M. Rae, A.E. Smith, J.S. Lilley and W.N. Catford, Scientific Program of the North-West Europe Nuclear Physics Conference, Edinburgh (1992), p.46.

18. “The Breakup of ^{24}Mg from an Oxygen Target”, C.D. Jones, D.L. Watson, G.J. Gyapong, R.M. Clark, B.R. Fulton, S.J. Bennett, M. Freer, J.T. Murgatroyd, N.M. Clarke, M. Leddy, W.D.M. Rae, A.E. Smith and W.N. Catford, Scientific Program of the North-West Europe Nuclear Physics Conference, Edinburgh (1992), p.47.

19. “Reaction Mechanism for the Breakup of ^{24}Mg”, G.J. Gyapong, C.D. Jones, D.L. Watson, S.J. Bennett, N.M. Clarke, B.R. Fulton, M. Leddy, J.T. Murgatroyd, W.D.M. Rae, P. Simmons and W.N. Catford, Scientific Program of the North-West Europe Nuclear Physics Conference, Edinburgh (1992), p.48.

20. “Hyperdeformed states of ^{36}Ar and ^{48}Cr in the cranked cluster model”, W.D.M.Rae and A.C.Merchant, Int. Nuclear Physics Conference, Wiesbaden, July 1992, Abstracts, No.1.2.50, ed. U.Grundinger.

21. “The Breakup of ^{24}Mg from an ^{16}O Target”, C.D. Jones, D.L. Watson, G.J. Gyapong, R.M. Clark, B.R. Fulton, S.J. Bennett, M. Freer, J.T. Murgatroyd, N.M. Clarke, M. Leddy, W.D.M. Rae, A.E. Smith and W.N. Catford, Int. Nuclear Physics Conference, Wiesbaden, July 1992, Abstracts, No.1.4.46, ed. U.Grundinger.

22. “Reaction Mechanism for the Breakup of ^{24}Mg”, G.J. Gyapong, C.D. Jones, D.L. Watson, S.J. Bennett, N.M. Clarke, B.R. Fulton, M. Leddy, J.T. Murgatroyd, W.D.M. Rae, P. Simmons and W.N. Catford, Int. Nuclear Physics Conference, Wiesbaden, July 1992, Abstracts, No.1.4.47, ed. U.Grundinger.

23. “^{16}O + ^{8}Be Breakup of ^{24}Mg from the ^{12}C(^{20}Ne,^{16}O^{8}Be)^{8}Be Reaction”, J.S. Pople, N.M. Clarke, B.R. Fulton, M.J. Leddy, J.T. Murgatroyd, Y. Chan, R.G. Stokstad, W.D.M. Rae, W.N. Catford, S.P. Fox, G.J. Gyapong, D.L. Watson and S.J. Bennett, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993, Programme and Abstracts p.31.

24. “The Search for a 7–α Chain State in ^{28}Si”, P.M. Simmons, W.D.M. Rae, T. Davinson, S.P. Fox, G.J. Gyapong, C.D. Jones, D.L. Watson, N.M. Clarke, B.R. Fulton, M. Leddy, J.T. Murgatroyd, A. Murphy, J.S. Pople, W.N. Catford and J. Lilley, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993, Programme and Abstracts p.34.

25. “Nuclear Molecule Configurations in ^{48}Cr”, A. St.J Murphy, S.J. Bennett, N.M. Clarke, B.R. Fulton, M.J. Leddy, J.T. Murgatroyd, J.S. Pople, G.J. Gyapong, S.P. Fox, C.D. Jones, D.L. Watson, W.N. Catford, W.D.M. Rae and P. Simmons, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993, Programme and Abstracts p.35.

26. “Systematics of alpha chain states in 4N-nuclei,”

A.C.Merchant and W.D.M.Rae, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993, Programme and Abstracts p.36.

27. “An Investigation into the Reaction Mechanism Populating Cluster States in 24Mg,” S.P. Fox, D.L. Watson, C.D. Jones, B.R. Fulton, N.M. Clarke, M.J. Leddy, J.S. Pople, A. S. Murphy, W.D. Rae, P.M. Simmons, W.N. Catford and G.J. Gyapong, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993 Programme and Abstracts p.37.

28. “Forward–Backward Asymmetry in Hyperdeformed ^{24}Mg Production,” S.M. Singer, W.N. Catford, G.J. Gyapong, S.P. Fox, C.D. Jones, D.L. Watson, N.M. Clarke, B.R. Fulton, M. Leddy, A. Murphy, J. Pople, P.M. Simmons and W.D.M. Rae, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993, Programme and Abstracts p.64.

29. “Systematics of 2-dimensional alpha cluster configurations in 4N-nuclei from ^{12}C to ^{44}Ti”, J.Zhang and W.D.M.Rae, IOP Conference on Nuclear and Particle Physics, University of Glasgow, March 1993 Programme and Abstracts p.76.

30. “Geometry and Collectivity in the Bloch-Brink Alpha Cluster Model,” J. Zhang, A.C. Merchant and W.D.M. Rae, IOP Annual Congress, Brighton, April 1994, Abstracts p.12.

31. “Calculations of particle decay widths of hyper-deformed nuclei using prolate spherical coordinates,” A.C. Merchant and W.D.M. Rae, IOP Annual Congress, Brighton, April 1994, Abstracts p.12.

32. “Break-up of ^{28}Si and ^{32}S to ^{12}C+^{16}O and ^{16}O+^{16}O,”

N. Curtis, N.M. Clarke, B.R. Fulton, S.J. Hall, M.J. Leddy, J.S. Pople, G. Tungate, R.P. Ward, W.N. Catford, G.J. Gyapong, S.M. Singer, C.D. Jones and P.H. Regan, IOP Annual Congress, Brighton, April 1994, Abstracts p.12.

33. “^{16}O+^{8}Be Break-up of ^{24}Mg ,”

J.S. Pople, N.M. Clarke, B.R. Fulton, M.J. Leddy, Y. Chan, R.G. Stokstad, W.D.M. Rae, W.N. Catford, G.J. Gyapong, S.P. Fox, D.L. Watson, S.J. Bennett and J.T. Murgatroyd, IOP Annual Congress, Brighton, April 1994, Abstracts p.13.

34. “Systematics of some 3–dimensional α–cluster configurations in 4N nuclei from 16O to ^{44}Ti,” J. Zhang, W.D.M. Rae and A.C. Merchant, IOP Annual Congress, Brighton, April 1994, Abstracts p.30.

35. “Excitation machanism of the 7–α chain state in ^{28}Si,” P.M. Simmons, W.D.M. Rae, J. Zhang, T. Davinson, S.P. Fox, G.J. Gyapong, C.D. Jones, D.L. Watson, N.M. Clarke, B.R. Fulton, M. Freer, M.J. Leddy, J.T. Murgatroyd, A. St.J. Murphy, J. Pople, W.N. Catford and J.S. Lilley, IOP Annual Congress, Brighton, April 1994, Abstracts p.33.

36. “High resolution measurement of the ^{12}C breakup state widths in ^{24}Mg,” N. Curtis, N.M. Clarke, B.R. Fulton, S.J. Hall, M.J. Leddy, A. St. J. Murphy, J.S. Pople, R.P. Ward, W.N. Catford, G.J. Gyapong, S.M. Singer, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, W.D.M. Rae and P.M. Simmons, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 1.

37. “^{16}O+^{8}Be Break-up of ^{24}Mg ,”

J.S. Pople, N.M. Clarke, B.R. Fulton, M.J. Leddy, Y. Chan, R.G. Stokstad, W.D.M. Rae, W.N. Catford, G.J. Gyapong, S.P. Fox, D.L. Watson, S.J. Bennett and J.T. Murgatroyd, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 2.

38. “The ^{12}C(^{12}C,^{12}C^{∗})^{12}C^{∗} ‘6–α Chain State’ Resonance,”

S.P.G. Chappell, D.L. Watson, S.P. Fox, C.D. Jones, W.D.M. Rae, P.M. Simmons, M.Freer, S.J. Hall and R.P. Ward, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 3.

39. “Inelastic scattering of ^{12}C+^{12}C,”

C.D. Jones, D.L. Watson, S.P.G. Chappell, S.P. Fox, P. Lee, M. Freer, S.J.A. Hall, G. Kelly, R. Le Marachel, R.P. Ward, W.N. Catford and S.M. Singer, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 4.

40. “A Study of the ^{12}C(^{16}O,^{12}C(0_{2}^{+}))^{16}O reaction,”

S.P. Fox, D.L. Watson, S.P.G. Chappell, C.D. Jones, P. Lee, M. Freer, S.J.A. Hall, G. Kelly, R. Le Marachel, R.P. Ward, W.N. Catford, G.J. Gyapong, S.M. Singer and W.D.M. Rae, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 10.

41. “Possible 2–dimensional α–cluster structures in ^{28}Si,”

G.J. Gyapong, W.N. Catford, S.M. Singer, S.P.G. Chappell, S.P. Fox, C.D. Jones, P. Lee, D.L. Watson, M. Freer, R. Le Marachel, G. Kelly and R.P. Ward, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 11.

42. “The search for a 7–α chain state in ^{28}Si,”

P.M. Simmons, W.D.M. Rae, J. Zhang, T. Davinson, S.P. Fox, G.J. Gyapong, C.D. Jones, D.L Watson, N.M. Clarke, B.R. Fulton, M. Freer, M.J. Leddy, J.T. Murgatroyd, A. St.J. Murphy, J.S. Pople, W.N. Catford and J.S. Lilley, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 13.

43. “Numerical algorithms for accurate angular momentum projection,”

J. Zhang and W.D.M. Rae, Cluster ’94, 6th Int. conf. on clusters in nuclear structure and dynamics, Strasbourg, France, Sep. 1994, Contributed paper 50.

44. “A Partial wave analysis of the shape eigenstate resonance,”

A.C. Merchant and W.D.M. Rae, IOP Annual Congress, Telford, March 1995, Abstract Book RNA3.3, p.104.

45. “Angular distribution measurements of the symmetric fission of ^{24}Mg,” S.M. Singer, W.N. Catford, G.J. Gyapong, S.P.G. Chappell, S.P. Fox, C.D. Jones, P. Lee, D.L. Watson, N.M. Clarke, M. Freer, B.R. Fulton, G.R. Kelly, M. Leddy, R.A. Le Marechal, A. Murphy, J.S. Pople, R.P. Ward, W.D.M. Rae and P.M. Simmons, IOP Annual Congress, Telford, March 1995, Abstract Book RNA4.1, p.105.

46. “Association of the ^{12}C+ ^{12}C breakup states in ^{24}Mg with the quasi-molecular resonances,” N. Curtis, N.M. Clarke, B.R. Fulton, S.J. Hall, M.J. Leddy, A. St. J. Murphy, J.S. Pople, R.P. Ward, W.N. Catford, G.J. Gyapong, S.M. Singer, S.P.G. Chappell, S.P. Fox, C.D. Jones, D.L. Watson, W.D.M. Rae and P.M. Simmons, IOP Annual Congress, Telford, March 1995, Abstract Book RNA4.2, p.105.

47. “B(E2) values in light self-conjugate nuclei,”

J. Zhang and W.D.M. Rae, IOP Annual Congress, Telford, March 1995, Abstract Book RNA4.3, p.106.

48. “Neutron angular distributions from the breakup of ^{19}C,”

J.C. Angélique, L. Axelsson, G. Bizard, W.N. Catford, N.M. Clarke, G. Costa, M. Freer, S. Grevy, D. Guillemaud-Mueller, G.J. Gyapong, F. Hanappe, B. Heusch, B. Jonson, C. Le Brun, F.R. Lecolley, F. Lefebvres, M. Lewitowicz, E. Liegard, F.M. Marqués, G. Martinez, A.C. Mueller, T. Nilsson, A. Ninane, G. Nyman, N.A. Orr, B. Peterson, F. Pougheon, K. Riisager, M.G. St. Laurent, Y. Schutz, M. Smedberg, O. Sorlin, L. Stuttge and D.D. Warner, Proc. Int. Conference on Exotic Nuclei and Atomic Masses, Arles, France, June 1995, ed. D. Guillemaud-Mueller, (Editions Frontières, 1995).

49. “MEGHA: A new detector array for charged particle spectroscopy,”

D.L. Watson for the Charissa Collaboration, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper A14.

50. “The ^{16}O* + ^{8}Be Mass Partition in the ^{24}Mg* LCC,”

P. Lee, C.D. Jones, S.P. Fox, S.P.G. Chappell, D.L. Watson, B.R. Fulton, N.M. Clarke, M. Freer, R.P. Ward, S.J. Hall, N. Curtis, R.A. Le Marechal, G.R. Kelly, W.D.M. Rae, P.M. Simmons, W.N. Catford, G.J. Gyapong and S.M. Singer, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper P24.

51. “Testing of two types of 2-dimensional position sensitive silicon detectors,” R.L. Cowin, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper P12.

52. “A coupled channels analysis of the resonances observed in ^{12}C + ^{12}C inelastic scattering at E_{c.m.} ∼ 33 MeV,” P.E. Fry, W.D.M. Rae and A.C. Merchant, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper P17.

53. “Study of ‘Six-alpha chain state’ resonance,”

R.A. Le Marechal, N.M. Clarke, M. Freer, B.R. Fulton, S. Hoad, C.D. Jones, G.R. Kelly, P. Lee, R.P. Ward and D.L. Watson, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper E18.

54. “Succesive alpha decay from excited states in ^{28}Si,”

G.R. Kelly, N.M. Clarke, M. Freer, B.R. Fulton, S. Hoad, R.A. Le Marechal and R.P. Ward, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper E19.

55. “Angular correlations for α-particle decay in the reaction^{12}C(^{12}C,^{12}C(3_{1}^{−}→^{8}Be(0_{1}^{+}) +α)^{12}C* near E_{c.m.} = 32.5 MeV,” S.P.G. Chappell and W.D.M. Rae, North-West Europe Nuclear Physics Conference, Amsterdam, April 1996, Contibuted Paper E20.

56. “Target dependence in breakup experiments,”

N. Curtis, M. Shawcross, W.N. Catford, B.R. Fulton, S.J. Hall, J.T. Murgatroyd, S.P.G. Chappell, R. Cowin, G. Dillon and D.L. Watson, IOP Conference on Nuclear Physics, University of York, 7-9th April 1997, contributed paper.

57. “Reaction mechanism populating states in ^{28}Si with pronounced ^{12}C + ^{16}O cluster structure,” M. Shawcross, N. Curtis, W.N. Catford, B.R. Fulton, J.T. Murgatroyd, S.J. Hall, S.P.G. Chappell, D.L. Watson, G. Dillon and R. Cowin, IOP Conference on Nuclear Physics, University of York, 7-9th April 1997, contributed paper.

58. “An angular correlation technique for investigating the reaction mechanism of 12C(^{24}Mg,^{12}C ^{12}C)^{12}C,” C.D. Freeman, B.R. Fulton, S.P.G. Chappell, N.M. Clarke, R. Cowin, R. Cunningham, G. Dillon, M. Freer, R. Hunt, S.J. Hall, G. Kelly, J.T. Murgatroyd, V. Pucknell, R. Ward and D.L. Watson, IOP Conference on Nuclear Physics, University of York 7-9th April 1997, contributed paper 2B1.

59. “Preliminary results from experiments using the Charissa PSSD array,” C.A. Bremner, S.P.G. Chappell and W.D.M. Rae, IOP Conference on Nuclear Physics, University of Liverpool 15-17th April 1997, contributed paper P16.

60. “Investigations of (^{16}O,^{12}C) reactions using the MEGHA strip array,” G.K. Dillon and S.P.G. Chappell, Nuclear Structure ’98 conference, Gatlingburg 10-15th August 1998, contributed paper.

61. “Structure of ^{11}Be from the (p,d) reaction in inverse kinematics,”

J.S. Winfield, S. Fortier, S. Pita, W.N. Catford, N.A. Orr, Y. Blumenfeld, R. Chapman, S.P.G. Chappell, N.M. Clarke, N. Curtis, M. Freer, S. Gales, K.L. Jones, H. Langevin-Joliot, H. Laurent, I.Lhenry, J.M. Maison, P. Roussel-Chomaz, M. Shawcross, M. Smith, K. Spohr, T. Suomijarvi and A. de Vismes, Nuclear Structure at the Extremes, Lewes 17-19th June 1988, contributed paper.

62. “Exotic molecular and halo states in ^{12,14}Be,”

M. Freer, N.A. Orr, M. Labiche, F.M. Marques, J.C. Angelique, L. Axelsson, B. Benoit, U. Bergmann, M.J.G. Borge, W.N. Catford, S.P.G. Chappell, N.M. Clarke, G. Costa, N. Curtis, A. D’Arrigo, F. D’Oliviera, E. de Goes Brennard, O. Dorvaux, B.R. Fulton, G. Gardina, C. Gregori, S. Grevy, D. Guillemaud-Mueller, F. Hanappe, B. Heusch, B. Jonson, G. Kelly, C. Le Brun, S. Leenhardt, M. Lewitowicz, K. Markenroth, M. Motta, A.C. Meuller, J.T. Murgatroyd, T. Nilsson, A. Ninane, G. Nyman, I. Piqueras, K. Risager, M.G. Saint-Lauerent, F. Sarazin, S. Singer, O. Sorlin, L. Stuttge and D.L. Watson, 2nd International conference on exotic nuclei and atomic masses, Michigan 23-27th June 1998, contributed paper.

63. “Study of ^{11}Be structure via the p(^{11}Be,^{10}Be)d reaction,”

S. Fortier, J.S. Winfield, S. Pita, W.N. Catford, N.A. Orr, Y. Blumenfeld, R. Chapman, S.P.G. Chappell, N.M. Clarke, N. Curtis, M. Freer, S. Gales, K.L. Jones, H. Langevin-Joliot, H. Laurent, I.Lhenry, J.M. Maison, P. Roussel-Chomaz, M. Shawcross, M. Smith, K. Spohr, T. Suomijarvi and A. de Vismes, 2nd International conference on exotic nuclei and atomic masses, Michigan 23-27th June 1998, contributed paper.

64. “Structure of ^{11}Be from the (p,d) reaction,”

S. Pita, S. Fortier, J.S. Winfield, W.N. Catford, N.A. Orr, Y. Blumenfeld, R.

Chapman, S.P.G. Chappell, N.M. Clarke, N. Curtis, A. de Vismes, M. Freer, S. Gales, K.L. Jones, H. Langevin, H. Laurent, I.Lhenry, J.M. Maison, P. Roussel-Chomaz, M. Shawcross, M. Smith, K. Spohr and T. Suomijarvi, INCP Paris 1998, contributed paper Vol. 1, p.407

65. “Energy levels in the unbound nucleus ^{13}Be,”

K.L. Jones et al., IOP Conference on nuclear and particle physics, Salford, April 1999, contributed paper.

66. “Study of single proton halo nuclei using transfer reactions,”

R.C. Lemmon et al., IOP Conference on nuclear and particle physics, Salford, April 1999, contributed paper.

67. “Reactions with the weakly bound Borromean nucleus ^{9}Be: the relevance to radioactive beam studies,” R.J. Wooliscroft et al., IOP Conference on nuclear and particle physics, Salford, April 1999, contributed paper.

68. “Techniques in the analysis of breakup reactions,”

P.J. Leask et al., IOP Conference on nuclear and particle physics, Salford, April 1999, contributed paper.

69. “High spin resonances in ^{12}C + ^{12}C scattering,”

C.A Bremner, S.P.G. Chappell and W.D.M. Rae, 7th Intern. Conf. on Clustering Aspects of Nuclear Structure and Dynamics, Rab, Island of Rab, Croatia, 14–19 June 1999, contributed paper.

70. “Coupled-channels calculations for the scattering of deformed light nuclei,” S.P.G. Chappell, C.A Bremner and W.D.M. Rae, 7th Intern. Conf. on Clustering Aspects of Nuclear Structure and Dynamics, Rab, Island of Rab, Croatia, 14–19 June 1999, contributed paper.

71. “Invariant mass spectroscopy of the unbound nucleus ^{13}Be,”

K.L. Jones et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

72. “Search for molecular states ^{16}C,”

P.J. Leask et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

73. “Study of the reaction ^{9}Be(^{9}Be,^{11}Be^{∗}) at 70 MeV,”

D. Mahboub et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

74. “High spin states observed via the ^{12}C + ^{16}O breakup of ^{28}Si,”

N.I. Ashwood et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

75. “Measurements of the ^{12}C + ^{12}C partial decay widths of ^{24}Mg observed in the 12C(^{16}O,^{12}C ^{12}C)^{4}He reaction,” M.-P. Nicoli et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

76. “Dependence on the density function in the double-folded model for the ^{9}Be + 208Pb system,” R.J. Wooliscroft et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

77. “Study of the energy dependence of the nucleus-nucleus interaction ^{9}Be + ^{12}C system,” V.A. Ziman et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

78. “Coupled channels calculations for the scatering of deformed light heavy ions: A challenge to the standard approach,” I.G. Boztosun and W.D.M. Rae, IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

79. “Search for ^{6}He + ^{4}He clustering in ^{10}Be,”

B.J. Greenhalgh et al., IOP Conference on nuclear physics, Birmingham, 2000, contributed paper.

80. “Coupled-channels calculations for the scattering of deformed light heavy ion reactions: A challenge to the standard approach,” I. Boztosun and W.D.M. Rae, Proc. 7th Int. Conf. on Clustering aspects of nuclear structure and dynamics, Ed. M. Korolija, Z. Basrak and R. Caplar, World-Scientific (2000) 143.

8 Other Output

1. “Probing Nuclear Molecules with Polarised Beams,”

SERC Annual Report 1987-88, p77

2. “Search for the Nuclear Molecule,”

SERC Annual Report 1991-92

3. “Nuclear sausage forged in carbon–carbon collisions”, Physics World, August 1992 issue, p.23

4. “Are Atoms Composed of Molecules?,”

New Scientist, 6 April 1991 issue, p.21

5. “Triangular alpha cluster geometries and harmonic oscillator shell structure”, J.Zhang and W.D.M.Rae, Oxford Internal Report OUNP–93–08, 1993.

6. “Search for γ-ray transitions between ^{12}C+^{12}C cluster states in ^{24}Mg,” A. Elanique, F. Haas, R.M. Freeman, C. Beck, R. Nouicer, D.L. Watson, C. Jones, R. Cowin, P. Lee, Z. Basrak, Strasbourg internal report.

7. “All together now,”

New Scientist, 27 February 1999 issue, p.27

8. “Paving the way for nuclear polymers,”

Physics World, May 1999 issue, p.22

9. “Of dumbells and doughnuts,”

New Scientist, 1 May 1999 issue, p.35

10. “Heavy water and the wartime race for nculear energy,” Review in Contemporary Physics, Vol. 42, No. 2 (2001) p.125-127